# Characterizations of voting rules based on majority margins

### v2

@@ Abstract
In the context of voting with ranked ballots, an important class of voting rules is the class of margin-based rules (also called pairwise rules). A voting rule is margin-based if whenever two elections generate the same head-to-head margins of victory or loss between candidates, then the voting rule yields the same outcome in both elections. Although this is a mathematically natural invariance property to consider, whether it should be regarded as a normative axiom on voting rules is less clear. In this paper, we address this question for voting rules with any kind of output, whether a set of candidates, a ranking, a probability distribution, etc. We prove that a voting rule is margin-based if and only if it satisfies some axioms with clearer normative content. A key axiom is what we call Preferential Equality, stating that if two voters both rank a candidate $x$ immediately above a candidate $y$, then either voter switching to rank $y$ immediately above $x$ will have the same effect on the election outcome as if the other voter made the switch, so each voter’s preference for $y$ over $x$ is treated equally.
@@ 1 Introduction
To state the point more formally, fix infinite sets $\mathcal{V}$ and $\mathcal{X}$ of voters and candidates, respectively. A profile is a function $\mathbf{P}:V\to\mathcal{O}(X)$ where $V$ is a nonempty finite subset of $\mathcal{V}$, $X$ is a nonempty finite subset of $\mathcal{X}$, and $\mathcal{O}(X)$ is the set of all strict weak orders on $X$; [^2] Recall that a strict weak order on $X$ is a binary relation $P$ on $X$ that is asymmetric and negatively transitive (i.e., if $(a,b)\not\in P$ and $(b,c)\not\in P$, then $(a,c)\not\in P$), which implies that $P$ is transitive. let us write $V(\mathbf{P})$ for $V$ and $X(\mathbf{P})$ for $X$ [^3] Allowing the set of voters to vary across profiles will be essential for our arguments. By contrast, all of our arguments would go through in a setting in which every profile has the same set of candidates. Given a profile $\mathbf{P}$ and candidates $a,b\in X(\mathbf{P})$, we define functions #$\#_{\mathbf{P}}$ and $\mathcal{M}_{\mathbf{P}}$ from $X(\mathbf{P})^{2}$ to $\mathbb{Z}$ as follows:
#$\#_{\mathbf{P}}(a,b)$ $=$ |$|\{i\in V(\mathbf{P})\mid(a,b)\in\mathbf{P}(i)\}|$ (1)
$\mathcal{M}_{\mathbf{P}}(a,b)$ $=$ #$\#_{\mathbf{P}}(a,b)-\#_{\mathbf{P}}(b,a).$ (2)
@@
Definition 1.1 is a mathematically natural invariance property to consider, but whether it should be regarded as a normative axiom on voting rules is less clear. In this paper, we address this question by proving that a voting rule is margin-based if and only if it satisfies some axioms with clearer normative content. A key axiom is what we call Preferential Equality (Definition 2.1), stating that if two voters both rank a candidate $x$ immediately above a candidate $y$, then either voter switching to rank $y$ immediately above $x$ will have the same effect on the election outcome as if the other voter made the switch, so each voter’s preference for $y$ over $x$ is treated equally. This implies (Lemma 2.4) that if there are some voters who rank $x$ immediately above $y$, then for any partition of those voters into groups $I$ and $J$ of equal size, the voters in $I$ switching to rank $y$ immediately above $x$ will have the same effect on the election outcome as if the voters in $J$ made the switch, so both group’s preferences for $y$ over $x$ are treated equally.
@@
In this election, IRV elects $R$, because $M$ is eliminated in the first round due to having the fewest first-place votes, and then $R$ beats $D$ in the second round. Now consider the 3% of Democratic voters who rank $R$ immediately above $M$, as well as 3% of the 32% of Republican voters who also rank $R$ immediately above $M$. In particular, consider the effect of either one of these groups switching to rank $M$ immediately above $R$:
@@
Our second key axiom is what Saari (2003) calls Neutral Reversal. This axiom states that adding to any profile a pair of reversed linear orders does not change the election outcome. The idea is that two voters with fully reversed preferences balance each other out, so the election outcome is unchanged. Again it is easy to see that margin-based voting rules satisfy Neutral Reversal, while Plurality and IRV do not.
Our first main theorem (Theorem 2.9) is that for the domain of linear profiles, in which each voter submits a complete linear order of the candidates, a voting rule is margin-based if and only if it satisfies Preferential Equality and Neutral Reversal. To obtain the analogous result for the domain of all profiles, we add a weak assumption of Homogeneity of the voting rule, as well as strengthening Preferential Equality to account for breaking ties in rankings (Theorem 3.8).
@@
where #$\#_{\mathbf{P}}$ is the function defined in (1) and $|V(\mathbf{P})|$ is the number of voters in $\mathbf{P}$. Thus, $\mathcal{H}(\mathbf{P})$ tells us not only how many voters prefer any $a$ to any $b$ but also the total number of voters, from which we can determine how many voters are indifferent between $a$ and $b$.
@@
A voting rule $F$ is a head-to-head voting rule if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $\mathcal{H}(\mathbf{P})=\mathcal{H}(\mathbf{P}^{\prime})$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. The rule $F$ is C2 if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if #$\#_{\mathbf{P}}=\#_{\mathbf{P}^{\prime}}$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. [^4] Although this was Fishburn’s (1977) original definition of C2, one could argue that since Fishburn’s paper was in the context of profiles of linear orders, his paper did not determine a unique extension of the C2 concept to profiles allowing ties. Nonetheless, we follow Brandt et al. 2018 in defining C2 for profiles with ties as in Definition 1.2.
@@
An example of a head-to-head voting rule that is neither margin-based nor even C2 is the Pareto rule that given a profile $\mathbf{P}$ outputs all the candidates who are not Pareto-dominated by any other candidate in $\mathbf{P}$. Whether one candidate Pareto-dominates another is not something we can tell just from $\mathcal{M}_{\mathbf{P}}$ or even #$\#_{\mathbf{P}}$, since we need to know the total number of voters to determine if every voter prefers candidate $x$ to candidate $y$ [^5] The Pareto rule falls into a class of rules that is intermediate between margin-based rules and the larger class of head-to-head rules, but incomparable with the class of C2 rules, namely the class of rules that can be computed using just the margins and the number of voters. We do not know of any literature discussing this class of rules, which could perhaps be called the quasi-margin-based rules. The alternative version of the Minimax rule defined using #$\#_{\mathbf{P}}$ in the main text below is not quasi-margin-based, though it is C2.
@@
We will also characterize when a head-to-head voting rule is margin-based. The normative importance of such characterizations largely has to do with questions about how to formulate common Condorcet voting rules when incomplete rankings (or ties in rankings) are allowed. Instead of measuring the strength of a majority preference for $x$ over $y$ in terms of the margin of victory of $x$ over $y$, some authors have considered measuring the strength of majority preference for $x$ over $y$ in other ways computable from #$\#_{\mathbf{P}}$ (see Schulze 2011, § 2.1). For example, although the Minimax voting rule (Simpson 1969, Kramer 1977) is now usually defined so as to output for a profile $\mathbf{P}$ the candidates $x$ who minimize the quantity $\mathrm{max}\{\mathcal{M}_{\mathbf{P}}(y,x)\mid y\in X(\mathbf{P})\}$, some have considered an alternative version that outputs the candidates $x$ who minimize $\mathrm{max}\{\#_{\mathbf{P}}(y,x)\mid y\in X(\mathbf{P})\}$. The latter “winning votes” version is C2 but not margin-based. To see how these definitions can output different sets of candidates, consider the profile at the top of Figure 1; below the profile we show both the margin graph of the profile, where an edge from candidate $x$ to candidate $y$ of weight $k$ indicates that $\mathcal{M}_{\mathbf{P}}(x,y)=k$, and the winning votes graph of the profile, where here an edge from candidate $x$ to $y$ of weight $k$ indicates that #$\#_{\mathbf{P}}(x,y)=k>\#_{\mathbf{P}}(y,x)$. The standard version of Minimax selects $c$ as the winner, whereas the winning votes version selects $a$ as the winner.
@@
Moreover, these distinctions are not merely academic. Examples from real political elections with ranked ballots are shown in Figure 2, where the standard version of Minimax selects Dornan, while the winning votes version selects Flanagan, and Figure 3, where the standard version selects Arab, while the winning votes version selects Worlobach; in each case, we only show candidates in the Smith set (Smith 1973), which is the smallest set of candidates such that every candidate in the set has a positive margin against every candidate outside the set. Table 1 shows the frequencies with which the two versions of Minimax disagree in four election databases, conditional on their being no Condorcet winner (both versions of Minimax elect the Condorcet winner whenever one exists).
@@
Figure 1: Above: a profile with ties. Below, left: the margin graph of the profile. Below, right: the winning votes graph of the profile.
2521 1657 1645 1051 768 681 486 381 190 180 $h$ $d$ $f$ $h$ $d$ $f$ $f$ $d$ $h$ $h$ $f$ $d$ $d$ $f$ $f$ $d$ $d$ $h$ $f$ $h$ $d$ $f$ $h$ $h$ $h$ $h$ $f$ $d$ $d$ $f$ DornanFlanaganHunter6026026026028686868621212121DornanFlanaganHunter299236543578357835783578
@@
Dataset # Profiles # Profiles with No Average Number Frequency of Different Winners
Condorcet Winner of Voters Conditional on No CW
Stable Voting Dataset 657 133 9.38 0.03
CIVS Dataset 22,477 6,050 26.74 0.09
PrefLib Politics (soi) 364 1 44,577.96 1.0
Otis 2022 dataset 458 2 74,840.35 0.5
Table 1: Frequencies with which Minimax based on margins and Minimax based on “winning votes” select different winners, conditional on there not being a Condorcet winner, in the elections from the Stable Voting Dataset 12-20-2024 (Holliday and Pacuit 2024), CIVS Dataset 2024-12-15 (Myers 2024a, b), PrefLib Politics (soi files as of 12-20-2024) (Mattei and Walsh 2013), and Otis 2022. Note that we restricted to elections that are not “test” elections in the CIVS Dataset.
All of this raises the question: what normative arguments are there for using the margin-based versions of voting rules? In the case of Minimax, previous work has identified reasons for favoring the margin-based version over the winning votes version: as Holliday and Pacuit (Forthcoming, Remark 3.11) observe, the former satisfies and the latter violates the axiom of Positive Involvement (stating that if a social choice correspondence selects a candidate $x$ in an initial profile $\mathbf{P}$, and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $x$ uniquely first, then $x$ should still be selected in $\mathbf{P}^{\prime}$). But what in general can be said about using margin-based rules as opposed to non-margin-based but still head-to-head rules? In what follows, we will provide answers to this question.
The rest of the paper is organized as follows. In Section 2, we characterize margin-based voting rules on the domain of linear profiles. In Section 3, we characterize which homogeneous voting rules on the domain of all profiles are margin-based. Finally, in Section 4, we characterize which head-to-head voting rules on the domain of all profiles are margin-based. We conclude in Section 5 with suggestions for future work. The Appendix gives analogues of theorems from Sections 3-4 when we allow noncomparability, as distinguished from indifference, in voters’ preference relations.

### v3

@@ Abstract
In the context of voting with ranked ballots, an important class of voting rules is the class of margin-based rules (also called pairwise rules). A voting rule is margin-based if whenever two elections generate the same head-to-head margins of victory or loss between candidates, the voting rule yields the same outcome in both elections. Although this is a mathematically natural invariance property to consider, whether it should be regarded as a normative axiom on voting rules is less clear. In this paper, we address this question for voting rules with any kind of output, whether a set of candidates, a ranking, a probability distribution, etc. We prove that a voting rule is margin-based if and only if it satisfies some axioms with clearer normative content. A key axiom is what we call Preferential Equality, stating that if two voters both rank a candidate $x$ immediately above a candidate $y$, then either voter switching to rank $y$ immediately above $x$ will have the same effect on the election outcome as if the other voter made the switch, so each voter’s preference for $y$ over $x$ is treated equally.
@@ 1 Introduction
To state the point more formally, fix infinite sets $\mathcal{V}$ and $\mathcal{X}$ of voters and candidates, respectively. A profile is a function $\mathbf{P}:V\to\mathcal{O}(X)$ where $V$ is a nonempty finite subset of $\mathcal{V}$, $X$ is a nonempty finite subset of $\mathcal{X}$ with $|X|\geq 2$, and $\mathcal{O}(X)$ is the set of all strict weak orders on XX; [^2] Recall that a strict weak order on $X$ is a binary relation $P$ on $X$ that is asymmetric and negatively transitive (i.e., if $(a,b)\not\in P$ and $(b,c)\not\in P$, then $(a,c)\not\in P$), which implies that $P$ is transitive. let us write $V(\mathbf{P})$ for $V$ and $X(\mathbf{P})$ for XX. [^3] Allowing the set of voters to vary across profiles will be essential for our arguments. By contrast, all of our arguments would go through in a setting in which every profile has the same set of candidates. Given a profile $\mathbf{P}$ and candidates $x,y\in X(\mathbf{P})$, we define functions $_{\mathbf{P}}$ and $\mathcal{M}_{\mathbf{P}}$ from $X(\mathbf{P})^{2}$ to $\mathbb{Z}$ as follows:
$\#_{\mathbf{P}}(x,y)$ $=$ $|\{i\in V(\mathbf{P})\mid(x,y)\in\mathbf{P}(i)\}|$ (1)
$\mathcal{M}_{\mathbf{P}}(x,y)$ $=$ $\#_{\mathbf{P}}(x,y)-\#_{\mathbf{P}}(y,x).$ (2)
@@
Definition 1.1 is a mathematically natural invariance property to consider, but whether it should be regarded as a normative axiom on voting rules is less clear. In this paper, we address this question by proving that a voting rule is margin-based if and only if it satisfies some axioms with clearer normative content. A key axiom is what we call Preferential Equality (Definition 2.1), stating that if two voters both rank a candidate $x$ immediately above a candidate $y$, then either voter switching to rank $y$ immediately above $x$ will have the same effect on the election outcome as if the other voter made the switch, so each voter’s preference for $y$ over $x$ is treated equally. This implies (Lemma 2.4) that if there are some voters who rank $x$ immediately above $y$, then for any partition of those voters into groups $I$ and $J$ of equal size, the voters in $I$ switching to rank $y$ immediately above $x$ will have the same effect on the election outcome as if the voters in $J$ made the switch, so both groups’ preferences for $y$ over $x$ are treated equally.
@@
In this election, IRV elects $R$, because $M$ is eliminated in the first round due to having the fewest first-place votes, and then $R$ beats $D$ in the second round. Now consider the 3% of Democratic voters who rank $R$ immediately above $M$, as well as 3% out of the 32% of Republican voters who also rank $R$ immediately above $M$. In particular, consider the effect of either one of these groups switching to rank $M$ immediately above $R$:
@@
There may be contexts in which treating voters equally does not require that a voting rule satisfy the axiom we have called ‘Preferential Equality’. For example, if the goal of the voting process is to maximize social utility, [^4] For arguments against taking preference intensities into account in collective choice, see Schwartz 1986, § 2.2. and the designer of the voting rule somehow knows that for any election to which the rule will be applied and any two voters, $i$ and $j$, the utility difference for $i$ between the candidate $i$ ranks 1st and the candidate $i$ ranks 2nd is greater than the utility difference for $j$ between the candidate $j$ ranks 2nd and the candidate $j$ ranks 3rd, [^5] This of course assumes the possibility of interpersonal comparisons of utility, contrary to Arrow’s (1963) view that “interpersonal comparison of utilities has no meaning and, in fact, that there is no meaning relevant to welfare comparisons in the measurability of individual utility” (p. 9). then one would not treat as equivalent $i$’s switching $x$ and $y$ in the 1st and 2nd positions of $i$’s ranking and $j$’s switching $x$ and $y$ in the 2nd and 3rd positions of $j$’s ranking. As another example, if one knows that voters are confident about their 1st choice but then flip a coin to determine who to rank 2nd vs. 3rd, one need not treat $i$’s switch and $j$’s switch as equivalent. However, in contexts without information about systematic differences in utility gaps or levels of noise between different positions in rankings, considerations of fairness and equal respect for voters motivate Preferential Equality.
Our second key axiom is what Saari (2003) calls Neutral Reversal. This axiom states that adding to any profile a pair of reversed linear orders does not change the election outcome. The idea is that two voters with fully reversed preferences balance each other out, so the election outcome is unchanged. This can be motivated by viewing an election as follows: by casting a ballot, a voter makes a contribution—independent of the contributions of other voters—to how well or poorly each candidate performs in the election; reversed rankings make opposite contributions; and the contribution of two voters is the sum of the contributions of each voter individually. Admittedly, one could instead adopt a holistic or super-additive view according to which, e.g., a pair of reversed rankings favors the middle-ranked candidates in the pair over the more “polarizing” candidates at the top and bottom of the respective rankings. [^6] Thanks to an anonymous referee for raising this point. For our purposes, the point is not that there is a decisive argument for Neutral Reversal but rather that there is a view motivating Neutral Reversal that does not presuppose that a voting method should be margin-based. The motivation for Neutral Reversal above is consistent with adopting voting methods such as Positive/Negative voting (García-Lapresta et al. 2010, Heckelman and Ragan 2021), where each voter awards $1$ point to their top-ranked candidate and 1-1 to their bottom-ranked candidate, or a hybrid between IRV and the Coombs method (Coombs 1964, Grofman and Feld 2004) that iteratively eliminates the candidate(s) with the lowest “score”, defined as their number of first-place votes minus their number of last-place votes. Neither of those methods is margin-based. On the other hand, it is easy to see that all margin-based voting rules satisfy Neutral Reversal, while Plurality and IRV do not.
Our first main theorem (Theorem 2.9) is that for the domain of linear profiles, in which each voter submits a complete linear order of the candidates, a voting rule is margin-based if and only if it satisfies Preferential Equality and Neutral Reversal. If we add a weak assumption of Homogeneity (Smith 1973) of the voting rule, we can replace Neutral Reversal with the less controversial axiom of Block Invariance (Holliday and Pacuit 2025), according to which the output of the voting rule remains unchanged if we add exactly one copy of each linear order of the candidates (Theorem 2.21). Homogeneity also allows us to characterize margin-based rules on the domain of all profiles, provided we strengthen Preferential Equality to account for breaking ties in rankings (Theorem 3.8).
@@
where $_{\mathbf{P}}$ is the function defined in (1) and $|V(\mathbf{P})|$ is the number of voters in $\mathbf{P}$. Thus, $\mathcal{H}(\mathbf{P})$ tells us not only how many voters prefer any $a$ to any $b$ but also the total number of voters, from which we can determine how many voters are indifferent between $a$ and $b$.
@@
A voting rule $F$ is a head-to-head voting rule if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $\mathcal{H}(\mathbf{P})=\mathcal{H}(\mathbf{P}^{\prime})$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. The rule $F$ is C2 if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $_{\mathbf{P}}=\#_{\mathbf{P}^{\prime}}$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. [^7] Although this was Fishburn’s (1977) original definition of C2 (except that, as in much of the literature, we do not exclude C1 methods from the class of C2 methods as Fishburn did), one could argue that since Fishburn’s paper was in the context of profiles of linear orders, his paper did not determine a unique extension of the C2 concept to profiles allowing ties.
@@
C2 voting rules are trivially head-to-head, and since the margins are computed from #\#, all margin-based rules are C2. An example of a C2 voting rule that is not margin-based is the weak Pareto rule that given a profile $\mathbf{P}$ outputs all the candidates who are not weakly Pareto-dominated by any other candidate in $\mathbf{P}$, where candidate $x$ weakly Pareto-dominates candidate $y$ if some voter prefers $x$ to $y$ and no voter prefers $y$ to $x$, i.e., $_{\mathbf{P}}(x,y)>0$ and $_{\mathbf{P}}(y,x)=0$. Whether one candidate weakly Pareto-dominates another is not something we can tell just from $\mathcal{M}_{\mathbf{P}}$, since, e.g., $\mathcal{M}_{\mathbf{P}}(x,y)=2$ is consistent with not only $_{\mathbf{P}}(x,y)=2$ and $_{\mathbf{P}}(y,x)=0$ but also $_{\mathbf{P}}(x,y)=4$ and $_{\mathbf{P}}(y,x)=2$. An example of a head-to-head voting rule that is neither margin-based nor even C2 is the strict Pareto rule that given a profile $\mathbf{P}$ outputs all the candidates who are not strictly Pareto-dominated by any other candidate in $\mathbf{P}$, where $x$ strictly Pareto-dominates $y$ if every voter prefers $x$ to $y$, i.e., $_{\mathbf{P}}(x,y)=|V(\mathbf{P})|$. Whether one candidate strictly Pareto-dominates another is not something we can tell just from $\mathcal{M}_{\mathbf{P}}$ or even $_{\mathbf{P}}$, since we need to know the total number of voters to determine if every voter strictly prefers $x$ to yy. [^8] The strict Pareto rule falls into a class of rules that is intermediate between margin-based rules and the larger class of head-to-head rules, but incomparable with the class of C2 rules and the class of C2w rules, namely the class of rules that can be computed using just the margins and the number of voters, since $x$ strictly Pareto-dominates $y$ if and only if $\mathcal{M}_{\mathbf{P}}(x,y)=|V(\mathbf{P})|$. We do not know of any literature discussing this class of rules, which could perhaps be called the quasi-margin-based rules. The weak Pareto rule and the alternative version of the Minimax rule defined using $_{\mathbf{P}}$ in the main text below are not quasi-margin-based, though they are C2. Of course, over the domain of linear profiles, the weak Pareto rule is equivalent to the strict Pareto rule and hence becomes C2.
Remark 1.3.
While we defined C2 in terms of $_{\mathbf{P}}$, which counts the number of voters who strictly prefer $x$ to $y$, one might consider instead counting the number of voters who weakly prefer $x$ to $y$. Where $^{w}_{\mathbf{P}}(x,y)=|\{i\in V(\mathbf{P})\mid(y,x)\not\in\mathbf{P}(i)\}|$, say that a voting rule $F$ is C2w if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $^{w}_{\mathbf{P}}=\#^{w}_{\mathbf{P}^{\prime}}$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. The notions of C2 and C2w are equivalent for voting rules defined on the domain of linear profiles, but they are not equivalent in general. All C2w voting rules are head-to-head since the $^{w}_{\mathbf{P}}$ function can be computed from the $_{\mathbf{P}}$ function together with $|V(\mathbf{P})|$ by the formula $^{w}_{\mathbf{P}}(x,y)=|V(\mathbf{P})|-\#_{\mathbf{P}}(y,x)$. In addition, all margin-based rules are C2w since $\mathcal{M}_{\mathbf{P}}(x,y)=\#^{w}_{\mathbf{P}}(x,y)-\#^{w}_{\mathbf{P}}(y,x)$. The above strict Pareto rule is C2w since $x$ strictly Pareto-dominates $y$ iff $^{w}_{\mathbf{P}}(y,x)=0$. Analogously, the weak Pareto rule is C2 but not C2w, as shown by the following two (anonymized) profiles having the same $^{w}$ counts:
11 11
$a$ $a,b$
$b$
22 11
$a$ $b$
$b$ $a$
Thus, the inclusion relations between head-to-head, C2, C2w, and margin-based rules can be summarized by the simple “diamond” diagram:
head-to-headC2C2wmargin-based
@@
We will also characterize when a head-to-head voting rule is margin-based, by way of characterizing C2 voting rules among head-to-head voting rules and characterizing margin-based voting rules among C2 voting rules. The normative importance of such characterizations largely has to do with questions about how to formulate common Condorcet voting rules when incomplete rankings (or ties in rankings) are allowed. Instead of measuring the strength of a majority preference for $x$ over $y$ in terms of the margin of victory of $x$ over $y$, some authors have considered measuring the strength of majority preference for $x$ over $y$ in other ways computable from $_{\mathbf{P}}$ (see Schulze 2011, § 2.1). For example, although the Minimax voting rule (Simpson 1969, Kramer 1977) is now usually defined so as to output for a profile $\mathbf{P}$ the candidates $x$ who minimize the quantity $\mathrm{max}\{\mathcal{M}_{\mathbf{P}}(y,x)\mid y\in X(\mathbf{P})\}$, some have considered an alternative version that outputs the candidates $x$ who minimize $\mathrm{max}\{\#_{\mathbf{P}}(y,x)\mid y\in X(\mathbf{P}),\,\#_{\mathbf{P}}(y,x)>\#_{\mathbf{P}}(x,y)\}$. The latter “winning votes” version is C2 but not margin-based (or C2w). To see how these definitions can output different sets of candidates, consider the profile at the top of Figure 1; below the profile we show both the margin graph of the profile, where an edge from candidate $x$ to candidate $y$ of weight $k$ indicates that $\mathcal{M}_{\mathbf{P}}(x,y)=k$, and the winning votes graph of the profile, where here an edge from candidate $x$ to $y$ of weight $k$ indicates that $_{\mathbf{P}}(x,y)=k>\#_{\mathbf{P}}(y,x)$. The standard version of Minimax selects $c$ as the winner, whereas the winning votes version selects $a$ as the winner.
@@
Moreover, these distinctions are not merely academic. Examples from real political elections with ranked ballots are shown in Figure 2, where the standard version of Minimax selects Dornan, while the winning votes version selects Flanagan, and Figure 3, where the standard version selects Arab, while the winning votes version selects Worlobach; in each case, we only show candidates in the Smith set (Smith 1973), which is the smallest set of candidates such that every candidate in the set has a positive margin against every candidate outside the set. Table 1 shows the frequencies with which the two versions of Minimax disagree in four election databases, conditional on there being no Condorcet winner (both versions of Minimax elect the Condorcet winner whenever one exists).
@@
Figure 1: Above: a profile with ties (three voters rank $a$ and $c$ as tied). Below, left: the margin graph of the profile. Below, right: the winning votes graph of the profile.
$2521$ $1645$ $1051$ $768$ $681$ $486$ $381$ $190$ $180$ $h$ $d$ $f$ $f$ $d$ $h$ $h$ $f$ $d$ dfd\ f $f$ $d$ dhd\ h fhf\ h $d$ $f$ $h$ $h$ $h$ $h$ $f$ $d$ $d$ $f$ DornanFlanaganHunter60260286862121DornanFlanaganHunter2992365435783578
@@
Dataset # Profiles # Profiles Avg. # Avg. # Frequency of Different Winners
with No CW of Voters of Cands. Conditional on No CW
Stable Voting Dataset 974 205 8.80 4.77 0.03
CIVS Dataset 22,477 6,050 26.74 9.77 0.09
PrefLib Politics (soi) 364 1 44,577.96 8.85 1.0
Otis 2022 dataset 458 2 74,840.35 6.10 0.5
Table 1: Frequencies with which Minimax based on margins and Minimax based on “winning votes” select different winners, conditional on there not being a Condorcet winner (CW), in the elections from the Stable Voting Dataset 1-28-2026 (Holliday and Pacuit 2026), CIVS Dataset 2024-12-15 (Myers 2024a, b), PrefLib Politics (soi files as of 12-20-2024) (Mattei and Walsh 2013), and Otis 2022. Note that we restricted to elections that are not “test” elections in the CIVS Dataset.
All of this raises the question: what normative arguments are there for using the margin-based versions of voting rules? In the case of Minimax, previous work has identified reasons for favoring the margin-based version over the winning votes version: as Holliday and Pacuit (2025, Remark 3.11) observe, the former satisfies and the latter violates the axiom of Positive Involvement (stating that if a social choice correspondence selects a candidate $x$ in an initial profile $\mathbf{P}$, and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one voter who ranks $x$ uniquely first, then $x$ should still be selected in $\mathbf{P}^{\prime}$). But what in general can be said about using margin-based rules as opposed to non-margin-based but still head-to-head rules? In what follows, we will provide answers to this question.
The rest of the paper is organized as follows. In Section 2, we characterize margin-based voting rules on the domain of linear profiles. In Section 3, we characterize which homogeneous voting rules on the domain of all profiles are margin-based. Finally, in Section 4, we characterize which head-to-head voting rules on the domain of all profiles are margin-based. We conclude in Section 5 with suggestions for future work. The Appendix gives analogues of theorems from Sections 3-4 when we allow noncomparability, as distinguished from indifference, in voters’ preference relations. Table 2 summarizes our main results using terminology to be defined.
result domain extra assumption axioms
Theorem 2.9 linear profiles none Preferential Equality
Neutral Reversal
Theorem 2.21 linear profiles Homogeneity Preferential Equality
Block Invariance
Theorem 3.8.1 LOBI Homogeneity Preferential Equality
Tiebreaking Compensation
Neutral Indifference
Theorem 3.8.2 all profiles Homogeneity Tiebreaking Compensation
Neutral Indifference
Theorem 4.4 all profiles head-to-head Nonlinear Neutral Reversal
Neutral Indifference
Table 2: Summary of axiomatic characterizations of margin-based voting rules. Over the indicated domain, a voting method that satisfies the extra assumption is margin-based if and only if it satisfies all of the indicated axioms. Two additional theorems appear in the Appendix.